OpenStudy (anonymous):

how long would it take an investment of $500 to double at a rate of 5.5% compounded quarterly

OpenStudy (anonymous):

\[i currently have 1/2(1+(0.055/4))^4xT\]

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 2\end{matrix}\right)(1+ \frac{ 0.055 }{ 4 })^{4timesT}\]

OpenStudy (perl):

solve $$ \Large 2\cdot 500 = 500(1+\frac{0.055}{4})^{4t} $$

Nnesha (nnesha):

haha latex >.< u got it for space type ~

OpenStudy (anonymous):

@perl ya thats how i got this equation i just dont know where to go from here.

OpenStudy (anonymous):

i need to solve for T

OpenStudy (anonymous):

the 1/2 would be on the other side of the equal sign wouldn't it?

OpenStudy (mathstudent55):

@perl 1/2 or 2 on the left side?

OpenStudy (perl):

woops, im thinking half-life

OpenStudy (perl):

$$ \Large{ 2\cdot 500 = 500(1+\frac{0.055}{4})^{4t} \\ \iff \\ 2= (1+\frac{0.055}{4})^{4t} } $$

OpenStudy (mathstudent55):

better

OpenStudy (anonymous):

ok now where from their?

OpenStudy (perl):

$$ \Large{ 2\cdot 500 = 500(1+\frac{0.055}{4})^{4t} \\ \iff \\ 2= (1+\frac{0.055}{4})^{4t} \\ \iff \\ \ln 2= \ln (1+\frac{0.055}{4})^{4t} \\ \iff \\ \ln 2= 4t \cdot \ln (1+\frac{0.055}{4}) } $$

OpenStudy (anonymous):

how do i get rid of ln? do they just cancel?

OpenStudy (perl):

you can't remove ln, but you can solve for t

OpenStudy (anonymous):

how so?

OpenStudy (perl):

$$ \Large{ 2\cdot 500 = 500(1+\frac{0.055}{4})^{4t} \\ \iff \\ 2= (1+\frac{0.055}{4})^{4t} \\ \iff \\ \ln 2= \ln (1+\frac{0.055}{4})^{4t} \\ \iff \\ \ln 2= 4t \cdot \ln (1+\frac{0.055}{4}) \\ \iff \\ \frac{\ln 2}{\ln (1+\frac{0.055}{4}) }= 4t } $$

OpenStudy (anonymous):

oh! alright i think i got it

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

12.689?

OpenStudy (anonymous):

@perl

OpenStudy (mathstudent55):

\(\Large \dfrac{\ln 2}{4\ln (1+\frac{0.055}{4}) }= t \) \(t = 12.689\)

OpenStudy (mathstudent55):

You are correct.

OpenStudy (perl):

yes

OpenStudy (anonymous):

sweet thanks a ton!